OpenStudy (anonymous):

What is the similarity ratio of two cubes with volumes of 64 cubic centimeters and 125 cubic centimeters? 1:5 2:3 4:5 16:25

OpenStudy (anonymous):

4:5

Directrix (directrix):

Any two cubes are similar. The cube of the scale factor is the ratio of the volumes. So, if the volumes are 64 and 125, then take the cube root of (64/125) = cube root of (4*4*4) / (5*5*5)= 4/5 or 4 to 5.